OpenStudy (anonymous):

If your solution is 0/3 does it means 0 or undefined?

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

0/3 = 0 since 0/x = 0 where x can be any number but 0

jimthompson5910 (jim_thompson5910):

it would be undefined if you had 3/0 since you CANNOT divide by zero

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}\frac{ 2x+1 }{ 3x^4-2 }= \frac{ x^4(\frac{ 2 }{ x^3 }+\frac{ 1 }{ x^4 }) }{ x^4(3-\frac{ 2 }{ x^4 }) }= \frac{ 0 }{ 3-0 }=?\]

OpenStudy (anonymous):

okay! got it thanks!(:

jimthompson5910 (jim_thompson5910):

also, in general, when you have a rational expression where the denominator has a larger degree than the numerator you are going to have an infinite limit of 0 since the denominator will always grow faster than the numerator

jimthompson5910 (jim_thompson5910):

but doing it this way also works

OpenStudy (anonymous):

okay! infinite limits of 0 is still 0 right so it shd be fine haha

jimthompson5910 (jim_thompson5910):

exactly, sounds like you got the hang of all this

OpenStudy (anonymous):

haha yup! btw one more question so does that means it has a horizonatal asymptote at 0 or it doent have?

jimthompson5910 (jim_thompson5910):

whenever the degree of the denominator is larger than the degree of the numerator, the infinite limit is always 0 and there's a horizontal asymptote at y = 0